You are deciding how to invest a total of $20,000 in two funds paying 5.5% or 7.5% simple interest. You want to earn a total of $1300 in interest from the investments each year. How much should you invest at 5.5% and how much should you invest at 7,5% to earn $1300?

Answer by mananth(16946)

Part I 5.50% per annum ---x *
Part II 7.50% per annum ---y
Total investment
x + 1 y= 20000 -------------1
Interest on both investments
5.50% x + 7.50% y= 1300
Multiply by 100
5.5 x + 7.5 y= 130000.00 --------2
Multiply (1) by -5.5
we get
-5.5 x -5.5 y= -110000.00
Add this to (2)
0 x 2 y= 20000
divide by 2
y = 10000 - x)
Part I 5.50% $ 10000
Part II 7.50% $ 10000

CHECK
10000 --------- 5.50% ------- 550.00
10000 ------------- 7.50% ------- 750.00
Total -------------------- 1300.00

m.ananth@hotmail.ca